OpenStudy (gabylovesyou):

The equation of line LM is y = 5x + 4. Write an equation of a line perpendicular to line LM in slope-intercept form that contains point (−3, 2).

OpenStudy (anonymous):

You know the slope of the new line has to have a slope that is the negative reciprocal of LM. start with that

OpenStudy (ranga):

1) What is slope of the line y = 5x + 4? 2) What is the slope of the line perpendicular to the line in step 1? 3) The equation of a line with slope m and passing through the point (x1, y1) is: y - y1 = m(x - x1). Simplify and put it in the form y = mx + b.

OpenStudy (gabylovesyou):

how do i find the slope

OpenStudy (anonymous):

It's the negative reciprocal of the slope of the original. The reciprocal is the inverse of a number. like so: 8 -> -1/8

OpenStudy (gabylovesyou):

ok.. i have to find the slope of what

OpenStudy (anonymous):

The original line. what is the slope of y=5x+4?

OpenStudy (gabylovesyou):

how do i find that

OpenStudy (ranga):

Compare it to y = mx + b where m is the slope.

OpenStudy (gabylovesyou):

slope is 5

OpenStudy (anonymous):

What is the negative reciprocal of 5?

OpenStudy (gabylovesyou):

-1/5

OpenStudy (gabylovesyou):

@cleats

OpenStudy (anonymous):

correct. now we have the following:\[\huge y - y_1 = m(x - x_1) \] we have m, we have x1 and y1 too

OpenStudy (gabylovesyou):

yes

OpenStudy (anonymous):

plug that stuff in and solve till you get y = mx + b

OpenStudy (gabylovesyou):

what is y ?

OpenStudy (gabylovesyou):

y1 is the slope ?

OpenStudy (gabylovesyou):

my internet is really bad... thats why i take forever to respond im so sorry

OpenStudy (anonymous):

m is the standard symbol for slope at least in lines. \[\huge y - y_1 = m(x - x_1)\](x1,y1) = (−3, 2) m = -1/5

OpenStudy (gabylovesyou):

y - 2 = -1/5(x - -3)

OpenStudy (anonymous):

yup. start by simplifying the parentheses. then distributing the -1/5

OpenStudy (gabylovesyou):

we have no x so how do i simplify parenthesis ?

OpenStudy (anonymous):

what does minus a -3 mean?

OpenStudy (gabylovesyou):

ummm

OpenStudy (anonymous):

lol i'll just write it out (x- -3)= x+3

OpenStudy (gabylovesyou):

oh lol ok ... so it would be -1/3 * x + 3

OpenStudy (gabylovesyou):

i mean -1/5 * x + 3

OpenStudy (anonymous):

y - 2 = -1/5(x + 3) yeah. distribute the -1/5

OpenStudy (gabylovesyou):

y - 2 = -1/5x + -1/5 + 3 ?

OpenStudy (anonymous):

y - 2 = -1/5x + -1/5 * 3 fixed

OpenStudy (gabylovesyou):

ok i do -1/5 * 3 ?

OpenStudy (anonymous):

yeah. -3/5

OpenStudy (gabylovesyou):

ok so we have y - 2 = -1/5x + -3/5

OpenStudy (anonymous):

add 2 to both sides aaaand we get?

OpenStudy (gabylovesyou):

y = -1/5x + 7/5

OpenStudy (anonymous):

yep. that's the answer

OpenStudy (gabylovesyou):

thank you so much!

OpenStudy (anonymous):

np, np